# Full-Duplex Millimeter Wave MIMO Channel Estimation: A Neural Network Approach

### v1

@@ III-A LS channel estimator
$\mathbf{\hat{H}}_{q,\text{LS}}=\frac{1}{\tau\sqrt{SNR_{q}}}\mathbf{\hat{Y}}_{q},$ (19)
@@ III-B MMSE channel estimator
$\begin{split}vec(\mathbf{Y_{\text{SI}}})=&\sqrt{SNR_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}vec(\mathbf{H_{\text{SI}}})+\\ &\sqrt{SNR_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}vec(\mathbf{H_{\text{UE}}})+vec(\mathbf{N}),\end{split}$ (20)
$\begin{split}vec(\mathbf{Y_{\text{UE}}})=&\sqrt{SNR_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}vec(\mathbf{H_{\text{UE}}})+\\ &\sqrt{SNR_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}vec(\mathbf{E_{\text{SI}}})+vec(\mathbf{N}),\end{split}$ (21)
@@
$vec(\mathbf{\hat{H}}_{\text{SI,MMSE}})=\sqrt{SNR_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}\left(SNR_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}+SNR_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}vec(\mathbf{Y_{\text{SI}}}).$ (27)
$vec(\mathbf{\hat{H}}_{\text{UE,MMSE}})=\sqrt{SNR_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}\left(SNR_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+SNR_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{E}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}vec(\mathbf{Y_{\text{UE}}}).$ (28)
$\mathbf{R_{\text{E}}}=\mathbf{R_{\text{SI}}}-SNR_{\text{SI}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}\left(SNR_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}+SNR_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}.$ (29)
@@
$\begin{split}&vec(\mathbf{\hat{H}_{\text{SI,MMSE}}})=\\ &\sqrt{SNR_{\text{SI}}}\left(\mathbf{R_{\text{SI}}}^{-1}+\tau SNR_{\text{SI}}\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}^{H}vec(\mathbf{Y_{\text{SI}}}),\end{split}$ (30)
$\begin{split}&vec(\mathbf{\hat{H}_{\text{UE,MMSE}}})=\\ &\sqrt{SNR_{\text{UE}}}\left(\mathbf{R_{\text{SI}}}^{-1}+\tau SNR_{\text{UE}}\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}^{H}vec(\mathbf{Y_{\text{SI}}}).\end{split}$ (31)
@@ III-C NN channel estimator
It has been shown in various studies that the wireless channel with large antenna arrays is spatially correlated, see e.g., [35]. To exploit this spatial correlation for channel estimation, we employ CNNs, which are suitable for capturing the local dependency of features among data points. The input to the CNN is the correlated pilot signal, i.e., $\mathbf{\hat{Y}}_{q}$, and the output is the estimated channel, $\mathbf{\hat{H}}_{q,\text{NN}}$, $q\in\{\text{SI},\text{UE}\}$. In convolutional layers, we apply a $3\times 3$ window size sliding through the whole input features with a unit stride size. Different numbers of hidden layers are employed, where each layer applies 64 convolutional kernels to extract features from the successive windows of its input features. The effect of the number of hidden layers will be carefully examined in Section IV.
@@
For training the CNN, a dataset consisting of $M_{\text{tr}}$ samples is generated with $(\mathbf{\hat{Y}}_{q}^{(n)},\mathbf{H}_{q}^{(n)})$ representing the $n$-th sample. We employ supervised learning, where $\mathbf{H}_{q}$ is regarded as the label during training. We apply min-max scaling to scale the dataset in the range $(0,1)$. Such normalization is highly recommended for training machine learning models. We consider MSE as the loss function
@@ III-D RX-TX channel mapping
The mathematical relationship between $\mathbf{h_{\text{UE,TX}}}$ and $\mathbf{h_{\text{UE,RX}}}$ is not well-defined due to the random AoA and the effects of multi-path propagation. On the other hand, based on universal approximation theory [30], feedforward NNs are capable of approximating any continuous function. This theory suggests that, given enough computational resources and data, it is possible to build NN models that can accurately approximate any function. Therefore, universal approximation theory inspires us to employ NNs to map the channel from the downlink UEs to the RX array to the channel from the TX array to the downlink UEs. Furthermore, recent studies [36] have demonstrated the existence of a spatial mapping function that can effectively map the channel from one set of antenna arrays to another. To accomplish this, we utilize an FNN with varying numbers of hidden layers. The input for FNN is the channel from the downlink UEs to the RX array, while the output is the channel from the TX array to the downlink UEs. Therefore, for the training of FNN, we collect $M_{\text{tr}}$ samples of $\mathbf{h_{\text{UE,RX}}}$ as the input and $\mathbf{h_{\text{UE,TX}}}$ as the corresponding label. Again, we apply min-max normalization and MSE loss function for training. We create the following data samples as the input and label of FNN to work with real-valued tensors:
@@ IV Simulation results
We consider an operating frequency is $28$ GHz, corresponding to a wavelength of $\lambda=10.71$ mm. We set the number of TX and RX arrays to 16, with a distance of $10\lambda$ between them, and antenna spacing in the transmit and the receive arrays is $d=\frac{\lambda}{2}$. The total number of downlink and uplink UEs is set to $K=8$. The path loss parameters are based on experimental results from [37], where the path loss constant at the reference distance is $\Gamma=-72$ dB, the path loss exponent is $\eta=2.92$, and the shadow fading standard deviation is $\sigma_{\text{sf}}=8.7$ dB. Unless otherwise claimed, the number of multi-path components is $P=5$ and AS is $\theta_{\text{AS}}=60^{\circ}$. The SI Rician factor is $\mu=40$ dB and we assume that the propagation SI cancellation is $\epsilon_{\text{SI}}=-40$ dB.
The NN-based channel estimators are trained on Python $3.7.16$ and implemented using the Keras libraries with a TensorFlow backend in the Jupyter Notebook environment. We employ Adam optimizer with a batch size of $512$ to update the network parameters. A dataset of $50,000$ samples is collected based on the channel model and it is split into $20,000$ samples for training, $20,000$ samples for validation, and $10,000$ samples for testing. The validation data is used to ensure that the model does not simply memorize the training data but learns meaningful aspects of the data for effective prediction. [^1] Upon acceptance, we will publish our source code for reproducibility.
@@ IV-B SI and UE channel estimation for different pilot dimensions
The NMSE of SI channel estimation vs SNRSI and SNRUE are provided in Figs. 7 and 8 for different pilot dimensions, respectively. The results demonstrate that the NN-based estimator outperforms the LS and MMSE estimators across considered pilot dimensions. Moreover, from Fig. 8, it is evident that the NN-based approach is more resilient to interference from UEs when compared to the MMSE and LS estimation methods. Using fewer pilot dimensions leads to a significant reduction in the quality of channel estimates for all three approaches. However, the reduction in pilot dimensions can allow more resources to be allocated to transmitting payload data. The trade-off between the performance of channel estimates and the number of pilot dimensions used depends on the required accuracy threshold for channel estimates and the system data rate.
We have also generated similar plots for UE channel estimation, showing the NMSE with respect to SNRUE and SNRSI. The plots are presented in Figs. 9 and 10, respectively, with varying pilot dimensions. We employ the MMSE estimator to cancel out the SI signal from the received pilot signal in LS, MMSE, and NN-based channel estimators for UE channel estimation, specifically when the pilot dimensions are $\tau=N_{\text{t}}$ and $\tau=K$. Note that we also utilize the error covariance matrix for the MMSE estimator when estimating UE channels as denoted in (28). From the results, it is evident that the NN-based approach using shorter pilot dimensions like $\tau=N_{\text{t}}$ or $\tau=K$, outperforms LS channel estimation with a pilot dimension of $\tau=N_{\text{t}}+K$. Even the MMSE channel estimator with longer pilot dimensions (e.g., $\tau=N_{\text{t}}+K$), is unable to outperform the NN-based technique with shorter pilot dimensions (e.g., $\tau=N_{\text{t}}$). These findings clearly demonstrate the superiority of NN in accurately estimating wireless channels, particularly in scenarios with lower SNRs and higher interference.
@@
IV-C Complexity Analysis
Finally, we compare the computational complexity of the LS, MMSE, and NN channel estimators in terms of FLOPs. The results are summarized in Table I, where $\zeta$ represents the side length of the convolutional window, and $f_{l-1}$ and $f_{l}$ denote the numbers of input and output convolutional kernel of the $l$-th layer. A bar plot is shown in Fig. 14 for SI channel estimation and pilot dimension $\tau=N_{\text{t}}$ based on the CNN architecture used in the simulation, i.e., $\zeta=3$, $f_{0}=f_{L}=2$, and $f_{i}=64$ for the remaining layers ($i=1,2,\dots,L-1$). Note that we have ignored the computational complexity of the covariance matrix calculation for MMSE estimation, as the channel statistics do not typically change during several coherence blocks. For the same reasons, we also ignored the computational complexity of the matrices in the MMSE formula that depend on spatial correlation matrices, assuming they can be pre-computed and used until the channel statistics change substantially. According to table I, the number of FLOPs exhibits quadratic growth with the number of received antennas in MMSE estimation, while for LS and NN-based estimation, it is linear. Furthermore, CNNs with no or one hidden layer require fewer FLOPs than MMSE estimation, while increasing the number of hidden layers significantly increases the computational complexity. In particular, adding a hidden layer with $64$ convolutional kernels and a $3\times 3$ window size requires about $1.5\times 10^{8}$ more FLOPs for estimating a MIMO channel with $16$ antennas.
TABLE I: Number of FLOPs for SI and UE channel estimation using LS, MMSE, and NN estimators.

### v2

@@ III-A LS channel estimator
$\mathbf{\hat{H}}_{q,\text{LS}}=\frac{1}{\tau\sqrt{{\mathrm{SNR}_{\text{q}}}}}\mathbf{\hat{Y}}_{q},$ (16)
@@ III-B MMSE channel estimator
$\begin{split}vec(\mathbf{Y_{\text{SI}}})=&\sqrt{\mathrm{SNR}_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}vec(\mathbf{H_{\text{SI}}})+\\ &\sqrt{\mathrm{SNR}_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}vec(\mathbf{H_{\text{UE}}})+vec(\mathbf{N}),\end{split}$ (17)
$\begin{split}vec(\mathbf{Y_{\text{UE}}})=&\sqrt{\mathrm{SNR}_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}vec(\mathbf{H_{\text{UE}}})+\\ &\sqrt{\mathrm{SNR}_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}vec(\mathbf{E_{\text{SI}}})+vec(\mathbf{N}),\end{split}$ (18)
@@
$vec(\mathbf{\hat{H}}_{\text{SI,MMSE}})=\sqrt{\mathrm{SNR}_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}\left(\mathrm{SNR}_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}+\mathrm{SNR}_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}vec(\mathbf{Y_{\text{SI}}}).$ (24)
$vec(\mathbf{\hat{H}}_{\text{UE,MMSE}})=\sqrt{\mathrm{SNR}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}\left(\mathrm{SNR}_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathrm{SNR}_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{E}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}vec(\mathbf{Y_{\text{UE}}}).$ (25)
$\mathbf{R_{\text{E}}}=\mathbf{R_{\text{SI}}}-\mathrm{SNR}_{\text{SI}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}\left(\mathrm{SNR}_{\text{SI}}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}\mathbf{\tilde{X}_{\text{SI}}}^{H}+\mathrm{SNR}_{\text{UE}}\mathbf{\tilde{X}_{\text{UE}}}\mathbf{R_{\text{UE}}}\mathbf{\tilde{X}_{\text{UE}}}^{H}+\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}\mathbf{R_{\text{SI}}}.$ (26)
@@
$\begin{split}&vec(\mathbf{\hat{H}_{\text{SI,MMSE}}})=\\ &\sqrt{\mathrm{SNR}_{\text{SI}}}\left(\mathbf{R_{\text{SI}}}^{-1}+\tau\mathrm{SNR}_{\text{SI}}\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}^{H}vec(\mathbf{Y_{\text{SI}}}),\end{split}$ (27)
$\begin{split}&vec(\mathbf{\hat{H}_{\text{UE,MMSE}}})=\\ &\sqrt{\mathrm{SNR}_{\text{UE}}}\left(\mathbf{R_{\text{SI}}}^{-1}+\tau\mathrm{SNR}_{\text{UE}}\mathbf{I}_{N_{\text{t}}N_{\text{r}}}\right)^{-1}\mathbf{\tilde{X}_{\text{SI}}}^{H}vec(\mathbf{Y_{\text{SI}}}).\end{split}$ (28)
@@ III-C NN channel estimator
It has been shown in various studies that the wireless channel with large antenna arrays is spatially correlated, see e.g., [38]. To exploit this spatial correlation for channel estimation, we employ CNNs, which are suitable for capturing the local dependency of features among data points.
The overall architecture of CNN for SI and UE channel estimation is depicted in Fig. 3.
Figure 3: The CNN architecture employed for SI and UE channel estimation. we apply $n$ ($n=0,1,2,10$) convolutional hidden layer(s) each with a $3\times 3$ window size and 64 convolutional kernels.
The input to the CNN is the correlated pilot signal, i.e., $\mathbf{\hat{Y}}_{q}$, and the output is the estimated channel, $\mathbf{\hat{H}}_{q,\text{NN}}$, $q\in\{\text{SI},\text{UE}\}$. In convolutional layers, we apply a $3\times 3$ window size sliding through the whole input features with a unit stride size. Different numbers of hidden layers are employed, where each layer applies 64 convolutional kernels to extract features from the successive windows of its input features. The effect of the number of hidden layers will be carefully examined in Section IV.
@@
For training, a dataset consisting of $M_{\text{tr}}$ samples is generated, with $(\mathbf{\hat{Y}}_{q}^{(n)},\mathbf{H}_{q}^{(n)})$ representing the $n$-th sample. We employ supervised learning, where $\mathbf{H}_{q}$ is regarded as the label during training. To obtain the ground truth channel samples, $\mathbf{H}_{q}^{(n)}$, for SI and UE channels, we follow the channel model introduced in Section II. Each realization of data samples in the dataset is generated based on this channel model, where, in each realization, the small-scale fading coefficient, large-scale fading, and AoA/AoD take on a new realization based on their corresponding distributions. Specifically, we draw samples from $\mathcal{CN}(0,1)$ for every multi-path component for small-scale fading, $\mathcal{U}\left[-\frac{\theta_{\text{AS}}}{2},\frac{\theta_{\text{AS}}}{2}\right]$ for every multi-path component for AoA/AoD, and $\mathcal{LN}(0,\sigma_{sf})$ for shadow fading in large-scale fading. After generating channel realizations, the received pilot signal for SI and UE channel estimation, $\mathbf{\hat{Y}}_{q}^{(n)}$, is generated by adding Gaussian noise to every realization and multiplying by the pilot matrix corresponding to different pilot dimensions. We apply min-max scaling to scale the dataset in the range $(0,1)$. Such normalization is highly recommended for training machine learning models. We consider MSE as the loss function
@@ III-D RX-TX channel mapping
The mathematical relationship between $\mathbf{h_{\text{UE,TX}}}$ and $\mathbf{h_{\text{UE,RX}}}$ is not well-defined due to the random AoA and the effects of multi-path propagation. On the other hand, based on universal approximation theory [30], feedforward NNs are capable of approximating any continuous function. This theory suggests that, given enough computational resources and data, it is possible to build NN models that can accurately approximate any function. Therefore, universal approximation theory inspires us to employ NNs to map the channel from the downlink UEs to the RX array to the channel from the TX array to the downlink UEs. Furthermore, recent studies [29] have demonstrated the existence of a spatial mapping function that can effectively map the channel from one set of antenna arrays to another. To accomplish this, we utilize an FNN with varying numbers of hidden layers. The input for FNN is the channel from the downlink UEs to the RX array, while the output is the channel from the TX array to the downlink UEs. Therefore, for the training of FNN, we collect $M_{\text{tr}}$ samples of $\mathbf{h_{\text{UE,RX}}}$ as the input and $\mathbf{h_{\text{UE,TX}}}$ as the corresponding label. To generate data samples for RX-TX channel mapping, we first generate the ground truth channel realizations for $\mathbf{h_{\text{UE,RX}}}$, and their corresponding labels for $\mathbf{h_{\text{UE,TX}}}$ are generated by adding its corresponding delay to each multi-path component. Again, we apply min-max normalization and MSE loss function for training. We create the following data samples as the input and label of FNN to work with real-valued tensors:
@@ IV Simulation results
We consider an operating frequency is $28$ GHz, corresponding to a wavelength of $\lambda=10.71$ mm. We set the number of TX and RX arrays to 16, with a distance of $10\lambda$ between them, and antenna spacing in the transmit and the receive arrays is $d=\frac{\lambda}{2}$. The total number of downlink and uplink UEs is set to $K=8$ and we assume all UEs are uniformly distributed within 20 m2. The path loss parameters are based on experimental results from [39], where the path loss constant at the reference distance is $\Gamma=-72$ dB, the path loss exponent is $\eta=2.92$, and the shadow fading standard deviation is $\sigma_{\text{sf}}=8.7$ dB. Unless otherwise claimed, the number of multi-path components is $P=5$ and AS is $\theta_{\text{AS}}=60^{\circ}$. The SI Rician factor is $\kappa=40$ dB and we assume that the propagation SI cancellation is $\epsilon_{\text{SI}}=-40$ dB [12].
The NN-based channel estimators are trained on Python $3.7.16$ and implemented using the Keras libraries with a TensorFlow backend in the Jupyter Notebook environment. We employ Adam optimizer with a batch size of $512$ to update the network parameters. A dataset of $50,000$ samples is collected based on the channel model and it is split into $20,000$ samples for training, $20,000$ samples for validation, and $10,000$ samples for testing. The validation data is used to ensure that the model does not simply memorize the training data but learns meaningful aspects of the data for effective prediction.
@@ IV-A Analysing NN-based channel estimator
To delve deeper into the problem of channel estimation with low-resolution ADCs, we followed the study in [40]. We assume that ADCs apply $b$-bit uniform scalar quantization with a set of $12^{b}-$ 1 thresholds denoted as $\{\tau_{1},\ldots,\tau_{2^{b}-1}\}$. The thresholds of the uniform quantizer will be
$\tau_{l}=(-2^{b-1}+l)\Delta_{b},\;\;\;l\in L=\{1,\ldots,2^{b}-1\},$ (39)
where $\Delta_{b}$ is the step size.
The quantization function is independently applied to the real and imaginary components of the signal, with its definition as follows:
$Q_{b}(r)=\begin{cases}\tau_{l}-\frac{\Delta_{b}}{2}&\text{if }r\in(\tau_{l-1},\tau_{l}]\text{ with }l\in L\\ (2^{b}-1)\frac{\Delta_{b}}{2}&\text{if }r\in(\tau_{2^{b}-1},\tau_{2^{b}}]\end{cases}$ (40)
Therefore, the received pilot signal after b-bit ADC follows:
$\mathbf{Y_{\text{b-bit}}}=Q_{b}(\mathbf{Y_{\text{SI}}}),$ (41)
The quantization step size is usually chosen to minimize the quantization MSE, and the optimal values are listed in Table I for their corresponding bits. The quantization distortion, $\eta_{b}$ is defined as
$\eta_{b}\triangleq\mathbb{E}\left[\frac{||Q_{b}(r)-r||^{2}_{F}}{||r||^{2}_{F}}\right],$ (42)
TABLE I: Optimal uniform quantizer [40]
Resolution (b) Step size ($\Delta_{b}$) Distortion ($\eta_{b}$)
1-bit $\sqrt{\frac{8}{\pi}}$ $1-\frac{2}{\pi}$
2-bit 0.996 0.1188
3-bit 0.586 0.0374
4-bit 0.335 0.0115
Figure 7: NMSESI vs SNRSI with a varying number of hidden layers for a full-duplex BS with few-bit ADCs.
We conducted simulations for 2-bit and 3-bit ADCs, and similar results can be extrapolated for higher resolutions. From Fig. 7, we can observe that increasing the number of hidden layers improves estimation with 2-bit and 3-bit ADCs, similar to the results for the 1-bit case. Furthermore, as expected, the higher the resolution of ADCs, the better the NMSE of channel estimates.
@@
Therefore, the choice of the depth of NN architecture is subject to the channel condition and cannot be regarded as universally applicable. Nevertheless, this observation underscores the significance of channel statistics in determining the architecture of NN for the channel estimation problem.
Finally, Fig. 9 shows the NMSE of SI channel estimation versus $\kappa$ with different values of $\kappa$ during training ($\kappa_{train}$). We exclude the plots with 1, 2, and 10 hidden layers for enhanced clarity. As it can be seen from this figure, depending on the value of $\kappa_{train}$, the NMSE of SI channel estimates behaves differently. More specifically, when it is trained on low $\kappa_{train}$, the NMSESI increases as $\kappa$ increases and when it is trained on high $\kappa_{train}$, the NMSESI decreases as $\kappa$ increases. On the other hand, training on different values of $\kappa_{train}$ ranging from 20 to 100 dB, shows that higher values of $\kappa$ lead to higher NMSESI.
Figure 9: NMSESI vs $\kappa$ with different values of $\kappa$ during training ($\kappa_{train}$).
IV-B SI and UE channel estimation for different pilot dimensions
@@
The NMSE of SI channel estimation vs SNRSI and SNRUE are provided in Figs. 10 and 11 for different pilot dimensions, respectively. The results demonstrate that the NN-based estimator outperforms the LS and MMSE estimators across considered pilot dimensions. Moreover, from Fig. 11, it is evident that the NN-based approach is more resilient to interference from UEs when compared to the MMSE and LS estimation methods. Using fewer pilot dimensions leads to a significant reduction in the quality of channel estimates for all three approaches. However, the reduction in pilot dimensions can allow more resources to be allocated to transmitting payload data. The trade-off between the performance of channel estimates and the number of pilot dimensions used depends on the required accuracy threshold for channel estimates and the system data rate.
We have also generated similar plots for UE channel estimation, showing the NMSE with respect to SNRUE and SNRSI. The plots are presented in Figs. 12 and 13, respectively, with varying pilot dimensions. We employ the MMSE estimator to cancel out the SI signal from the received pilot signal in LS, MMSE, and NN-based channel estimators for UE channel estimation, specifically when the pilot dimensions are $\tau=N_{\text{t}}$ and $\tau=K$. Note that we also utilize the error covariance matrix for the MMSE estimator when estimating UE channels as denoted in (25). From the results, it is evident that the NN-based approach using shorter pilot dimensions like $\tau=N_{\text{t}}$ or $\tau=K$, outperforms LS channel estimation with a pilot dimension of $\tau=N_{\text{t}}+K$. Even the MMSE channel estimator with longer pilot dimensions (e.g., $\tau=N_{\text{t}}+K$), is unable to outperform the NN-based technique with shorter pilot dimensions (e.g., $\tau=N_{\text{t}}$). These findings clearly demonstrate the superiority of NN in accurately estimating wireless channels, particularly in scenarios with lower SNRs and higher interference.
@@
IV-C Distribution shift
TABLE II: DeepMIMO dataset parameters
DeepMIMO parameters Value
Number of paths 10
Active UEs from row 550 to 1650
Active BS number BS 3
Bandwidth 50 MHz
BS antennas $N_{x}$ = 1, $N_{y}$ = 16, $N_{z}$ = 1
UE antennas $N_{x}$ = 1, $N_{y}$ = 1, $N_{z}$ = 1
When a learning model is trained on a particular dataset with a specific distribution of features, it learns to make predictions based on the patterns inherent in that dataset. However, deploying such models in real-world scenarios often involves testing them with datasets that may differ in distribution from the training data. This misalignment between training and test distributions can lead to a phenomenon known as distribution shift, where the model’s performance deteriorates due to its inability to generalize effectively to unseen data and it poses significant challenges in testing the efficacy and reliability of learning models.
To analyze the effect of distribution shift on our trained models, we utilize the DeepMIMO dataset in the test phase. This dataset constructs the MIMO channels based on ray-tracing simulation from Remcom Wireless InSite [41]. Specifically, we use the ‘O1’ scenario at 28 GHz with the parameters specified in Table II. We generate 4977 channel realizations from the DeepMIMO dataset to test the NN model for UE channel estimation trained based on the channel data from geometrical channel models introduced in Section II-A. The same can be applied for SI channel estimation, but for brevity, we do not include the results for this case. As seen from Fig. 17, due to distribution mismatch, the NMSE degrades in the test phase. Furthermore, we observe that NNs with fewer hidden layers generalize better compared to deeper CNNs since the number of trainable parameters is much higher in deep CNNs. This underscores the benefits of utilizing simpler NN architectures in terms of generalization abilities.
Figure 17: Performance of NN channel estimator for UE channel estimation with distribution shift.
IV-D Complexity Analysis
Finally, we compare the computational complexity of the LS, MMSE, and NN channel estimators in terms of FLOPs. The results are summarized in Table III, where $\zeta$ represents the side length of the convolutional window, and $f_{l-1}$ and $f_{l}$ denote the numbers of input and output convolutional kernel of the $l$-th layer. A bar plot is shown in Fig. 18 for SI channel estimation and pilot dimension $\tau=N_{\text{t}}$ based on the CNN architecture used in the simulation, i.e., $\zeta=3$, $f_{0}=f_{L}=2$, and $f_{i}=64$ for the remaining layers ($i=1,2,\dots,L-1$). Note that we have ignored the computational complexity of the covariance matrix calculation for MMSE estimation, as the channel statistics do not typically change during several coherence blocks. For the same reasons, we also ignored the computational complexity of the matrices in the MMSE formula that depend on spatial correlation matrices, assuming they can be pre-computed and used until the channel statistics change substantially. According to table III, the number of FLOPs exhibits quadratic growth with the number of received antennas in MMSE estimation, while for LS and NN-based estimation, it is linear. Furthermore, CNNs with no or one hidden layer require fewer FLOPs than MMSE estimation, while increasing the number of hidden layers significantly increases the computational complexity. In particular, adding a hidden layer with $64$ convolutional kernels and a $3\times 3$ window size requires about $1.5\times 10^{8}$ more FLOPs for estimating a MIMO channel with $16$ antennas.
TABLE III: Number of FLOPs for SI and UE channel estimation using LS, MMSE, and NN estimators.